# Bridging Privacy and Robustness for Trustworthy Machine Learning

### v4

@@ Appendix B Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
Appendix C Generalized Relationship between LDP and MBP
In this section, we derive the generalized relationship between LDP and MBP.
C-A LDP implies MBP
The following lemma asserts that when the adversary’s prior belief distribution is close to uniform, the mapping function provides a high level of privacy protection by limiting the information leakage to a maximum of $\xi$ in a Bayesian sense when the adversary observes a transformed data point $w$.
Lemma C.1 (LDP implies MBP).
Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution. That is, $\forall d,d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$. If $f_{W|D}(\cdot|\cdot)$ is $\xi$-LDP. That is, the mapping $f_{W|D}(\cdot|\cdot)$ satisfies that
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
then $f_{W|D}(\cdot|\cdot)$ is a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy. That is, $\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|\leq\xi$, $\forall w\in\mathcal{W},d\in\mathcal{D}$.
@@ C-A LDP implies MBP
From the definition of $\xi$-LDP, we know that
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-\xi},e^{\xi}],$
for all $d,d^{\prime}\in\mathcal{D}$. Notice that
$e^{-\xi}\leq\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\leq e^{\xi}.$
It implies that $f_{D|W}(d|w)\leq f_{D|W}(d^{\prime}|w)\cdot e^{\xi}\cdot\frac{f_{D}(d)}{f_{D}(d^{\prime})}\leq f_{D|W}(d^{\prime}|w)\cdot e^{\xi+\epsilon}$.
We assume that $\frac{f_{D}(d)}{f_{D}(d^{\prime})}\in[e^{-\epsilon},e^{\epsilon}]$. Notice that
$\sum_{d\in\mathcal{D}}f_{D|W}(d|w)=1.$
For any fixed $\tilde{d}\in\mathcal{D}$, we have that
11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
$\leq f_{D|W}(\tilde{d}|w)+e^{\xi+\epsilon}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
$=(1+e^{\xi+\epsilon}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
Therefore, we have that
$f_{D|W}(\tilde{d}|w)\geq\frac{1}{1+e^{\xi+\epsilon}\cdot(|\mathcal{D}|-1)}.$
Similarly, we have that
11 11 $=\sum_{d\in\mathcal{D}}f_{D|W}(d|w)$
$=f_{D|W}(\tilde{d}|w)+\sum_{d\in\mathcal{D}\setminus\tilde{d}}f_{D|W}(d|w)$
$\geq f_{D|W}(\tilde{d}|w)+e^{-\xi-\epsilon}\cdot f_{D|W}(\tilde{d}|w)\cdot(|\mathcal{D}|-1)$
$=(1+e^{-\xi-\epsilon}\cdot(|\mathcal{D}|-1))\cdot f_{D|W}(\tilde{d}|w).$
Therefore, we have
$f_{D|W}(\tilde{d}|w)$ $\leq\frac{1}{1+e^{-\xi-\epsilon}\cdot(|\mathcal{D}|-1)}$
$\leq\frac{e^{\xi+\epsilon}}{|\mathcal{D}|}.$
Therefore, $\forall\tilde{d}\in\mathcal{D}$, we have that
$f_{D|W}(\tilde{d}|w)\leq e^{\xi+\epsilon}f_{D}(\tilde{d}).$
We say a mechanism is $(\xi+\epsilon)$-MBP if $\forall w\in\mathcal{W}^{d}$,
$\frac{f(d|w)}{f(d)}=\frac{f(w|d)f(d)}{f(w)f(d)}=\frac{f(w|d)}{f(w)}\leq e^{\xi+\epsilon}.$ (11)
Therefore, $f_{D|W}(\cdot|\cdot)$ is $(\xi+\epsilon)$-MBP. ∎
C-B MBP implies LDP
The following lemma establishes a connection between MBP and LDP under the assumption that the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution. It states that if a mapping $f_{W|D}(\cdot|\cdot)$ guarantees $\xi$-MBP, then it also satisfies $(2\xi+\epsilon)$-LDP. In other words, if the log-ratio of the posterior belief $f_{D|W}(d|w)$ and the prior belief $f_{D}(d)$ is bounded by $\xi$ for all $w\in\mathcal{W}$ and $d\in\mathcal{D}$, and the prior belief distribution is "$\epsilon$-close" to uniform (i.e., $|\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$ for all $d,d^{\prime}$), then the ratio of the conditional probabilities $f_{W|D}(w|d)$ and $f_{W|D}(w|d^{\prime})$ is bounded within the range $[e^{-2\xi-\epsilon},e^{2\xi+\epsilon}]$ for any $w\in\mathcal{W}$ and $d,d^{\prime}\in\mathcal{D}$. This lemma provides a useful connection between MBP and LDP, which can be leveraged to analyze the privacy guarantees of algorithms satisfying MBP under certain conditions on the prior belief distribution.
Lemma C.2 (MBP implies LDP).
Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution. That is, $\forall d,d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$. Let $f_{W|D}(\cdot|\cdot)$ be a privacy-preserving mapping that guarantees $\xi$-maximum Bayesian privacy. That is, $\xi=\max_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$. Then, the mapping $f_{W|D}(\cdot|\cdot)$ is $(2\xi+\epsilon)$-Local Differential Private:
$\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\in[e^{-2\xi},e^{2\xi}].$
Proof.
From the definition of maximum Bayesian privacy leakage, we know that
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\xi},e^{\xi}],$
for any $w\in\mathcal{W}$ and $d\in\mathcal{D}$. Therefore, for any $d,d^{\prime}\in\mathcal{D}$ we have
$e^{-2\xi-\epsilon}\leq\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\cdot e^{-\epsilon}\leq\frac{f_{W|D}(w|d)}{f_{W|D}(w|d^{\prime})}\leq\frac{f_{D|W}(d|w)f_{D}(d^{\prime})}{f_{D|W}(d^{\prime}|w)f_{D}(d)}\cdot e^{\epsilon}\leq e^{2\xi+\epsilon}.$
From the definition of Differential Privacy, we know that it is $(2\xi+\epsilon)$-Differentially Private. ∎
Combining the results in C-A and C-B, we are now ready to analyze the generalized relationship between MBP and LDP, which is shown in the following theorem.
Theorem C.3 (Relationship between MBP and LDP).
Consider the scenario where the prior belief distribution of the adversary is "$\epsilon$-close" to a uniform distribution. That is, $\forall d,d^{\prime},\ |\log\frac{f_{D}(d)}{f_{D}(d^{\prime})}|\leq\epsilon$. If an algorithm is $\xi$-LDP, then it is $(\xi+\epsilon)$-MBP. If an algorithm is $\xi$-maximum Bayesian privacy, then it is $(2\xi+\epsilon)$-LDP.
Appendix D Generalized Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
In this section, we introduce the generalized relationship between average Bayesian privacy and maximum Bayesian privacy.
Theorem D.1 (Relationship Between MBP and ABP).
Assume that the prior belief of the attacker satisfies that
$\frac{f^{\mathcal{B}}_{D}(d)}{f_{D}(d)}\in[e^{-\epsilon},e^{\epsilon}].$
Let MBP be defined using Definition IV.3. For any $\epsilon_{p,m}\geq 0$, let $f_{W|D}(\cdot|\cdot)$ be a privacy preserving mapping that guarantees $\epsilon_{p,m}$-maximum Bayesian privacy. That is,
$\frac{f_{D|W}(d|w)}{f_{D}(d)}\in[e^{-\epsilon_{p,m}},e^{\epsilon_{p,m}}],$
where $\epsilon_{p,a}$ is introduced in Definition IV.1.
Proof.
Let $F^{\mathcal{M}}=\frac{1}{2}(F^{\mathcal{A}}+F^{\mathcal{B}})$. We have
Therefore, we have that
Let $w^{\star}=\arg\max_{w\in\mathcal{W}}f_{D|W}(d|w)$. Then we have that
[…]
$\leq e^{\epsilon_{p,m}}\cdot f_{D}(d)$
$\leq e^{\epsilon_{p,m}+\epsilon}\cdot f^{\mathcal{B}}_{D}(d).$ (12)
[…]
$\geq e^{-\epsilon_{p,m}}\cdot f_{D}(d)$
$\geq e^{-(\epsilon_{p,m}+\epsilon)}\cdot f^{\mathcal{B}}_{D}(d).$ (13)
Combining Eq. (D) and Eq. (D), we have
Now we analyze according to the following two cases.
Case 1: $f^{\mathcal{B}}_{D}(d)\leq f^{\mathcal{A}}_{D}(d)$. Then we have
Case 2: $f^{\mathcal{B}}_{D}(d)\geq f^{\mathcal{A}}_{D}(d)$. Then we have
Therefore,
Combining Eq. (D), Eq. (D) and Eq. (14), we have that
Therefore,
Appendix E Analysis for Theorem IV.7
Theorem E.1 (PAC Robustness of $\epsilon$-MBP Protection Mechanisms).
Let $\mathcal{M}$ be an $(\epsilon,\beta,\alpha)$-MBP (Definition IV.5) algorithm. Then $\mathcal{M}$ is $\alpha$-PAC-Robust with probability $1-\gamma$, where $\gamma\in(0,1)$, $n$ represents the size of the dataset, and $\beta=\frac{\gamma}{1+4\epsilon}$ represents the error probability.
Proof.
$\mathcal{M}$ is an $(\epsilon,\beta,\alpha)$-MBP algorithm for the estimation of statistic $\mu$. We know that the protection mechanism $\mathcal{M}$ is $\epsilon$-MBP (measured using Definition IV.3). From the relationship between MBP and LDP (Theorem IV.4), we know that $\mathcal{M}$ is $2\epsilon$-LDP. From the definition of private algorithm, we know that for the original private information $d$, we have that
@@ Appendix D Generalized Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
From Definition IV.6, $\mathcal{M}$ is $\alpha$-PAC-Robust with probability $1-\gamma$.
Appendix F Practical Estimation for Maximum Bayesian Privacy
@@
When the extent of the attack is fixed, $\xi$ becomes a constant. Given Assumption G.1, we can infer that
@@
We then estimate $\hat{f}_{{D}|{W}}(d|w_{m})$ with $\{w_{m}\}_{m=1}^{M}$. As mentioned in subsection III-A, the attacker can conduct Bayesian inference attack, i.e., estimate dataset $d$ with $\hat{d}$ that optimizes posterior belief, denoted by:
@@ F-B Estimation Error of MBP
The following lemma states that the estimated value $\hat{\kappa}_{1}(d)$ is very likely to be close to the true value $\kappa_{1}(d)$. The probability of the estimate being within an error margin that is a small percentage ($\epsilon$ times) of the true value is very high—specifically, at least $1-2\exp(-\frac{\epsilon^{2}T\kappa_{1}(d)}{3})$. This means that the larger the number $T$ or the smaller the error parameter $\epsilon$, the more confident we are that our estimate is accurate. For the detailed analysis, please refer to Zhang et al. (2023e).
Lemma F.1.
@@
Appendix G Practical Estimation for Average Bayesian Privacy
In the scenario of federated learning, Zhang et al. (2023e) derived the upper bound of privacy leakage based on the specific form of $\text{TV}(P^{\mathcal{O}}||P^{\mathcal{D}})$ of $\mathcal{M}$ according to this equation: $\widetilde{\epsilon}_{p}=2C_{1}-C_{2}\cdot{\text{TV}}(P^{\mathcal{O}}||P^{\mathcal{D}})$. In this section, we elaborate on the methods we propose to estimate $C_{1}$ and $C_{2}$. First, we introduce the following assumption.
Assumption G.1.
@@ Appendix G Practical Estimation for Average Bayesian Privacy
G-A Estimation for $C_{1}=\sqrt{{\text{JS}}(F^{\mathcal{O}}||F^{\mathcal{B}})}$
@@
G-B Estimation for $C_{2}=\frac{1}{2}(e^{2\xi}-1)$
Recall that the Maximum Bayesian Privacy, is defined as $\xi=\max_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$ and represents the maximal privacy leakage across all possible information $w$ released. The quantity $\xi$ remains constant when the extent of the attack is fixed. Under Assumption G.1, we have that
@@ G-B Estimation for $C_{2}=\frac{1}{2}(e^{2\xi}-1)$
Ultimately, the computation of $C_{1}$ and $C_{2}$ hinges on the estimation of $\hat{f}_{{D}|{W}}(d|w_{m})$, which is further elaborated upon in Section F-A.
G-C Estimation Error of ABP
In this section, we analyze the estimation error of ABP. To facilitate the analysis of the error bound, we consider that both the data and the model parameter are in discrete form. For the detailed analysis, please refer to Zhang et al. (2023e).
@@ G-C Estimation Error of ABP
Theorem G.1 (The Estimation Error of $C_{1}$).

### v5

@@ Appendix C Relationship Between Average Bayesian Privacy and Maximum Bayesian Privacy
C.1 Theoretical Relationship between PAC Robustness and MBP
Theorem C.7 (PAC Robustness of $\epsilon$-MBP Protection Mechanisms).
Let $\mathcal{M}$ be an $(\epsilon,\beta,\alpha)$-MBP (Definition 4.5) algorithm. Then $\mathcal{M}$ is $\alpha$-PAC-Robust with probability $1-\gamma$, where $\gamma\in(0,1)$, $n$ represents the size of the dataset, and $\beta=\frac{\gamma}{1+4\epsilon}$ represents the error probability.
@@ C.1 Theoretical Relationship between PAC Robustness and MBP
$\mathcal{M}$ is an $(\epsilon,\beta,\alpha)$-MBP algorithm for the estimation of statistic $\mu$. We know that the protection mechanism $\mathcal{M}$ is $\epsilon$-MBP (measured using Definition 4.3). From the relationship between MBP and LDP (Lemma 4.1), we know that $\mathcal{M}$ is $2\epsilon$-LDP. From the definition of private algorithm, we know that for the original private information $d$, we have that
@@
From Definition 4.6, $\mathcal{M}$ is $\alpha$-PAC-Robust with probability $1-\gamma$.
Appendix D Practical Estimation for Maximum Bayesian Privacy
@@
When the extent of the attack is fixed, $\xi$ becomes a constant. Given Assumption E.1, we can infer that
@@
We then estimate $\hat{f}_{{D}|{W}}(d|w_{m})$ with $\{w_{m}\}_{m=1}^{M}$. Recall that the attacker can conduct Bayesian inference attack, i.e., estimate dataset $d$ with $\hat{d}$ that optimizes posterior belief, denoted by:
@@ D.2 Estimation Error of MBP
The following lemma states that the estimated value $\hat{\kappa}_{1}(d)$ is very likely to be close to the true value $\kappa_{1}(d)$. The probability of the estimate being within an error margin that is a small percentage ($\epsilon$ times) of the true value is very high—specifically, at least $1-2\exp(-\frac{\epsilon^{2}T\kappa_{1}(d)}{3})$. This means that the larger the number $T$ or the smaller the error parameter $\epsilon$, the more confident we are that our estimate is accurate. For the detailed analysis, please refer to [64].
Lemma D.8.
@@
Appendix E Practical Estimation for Average Bayesian Privacy
In the scenario of federated learning, [64] derived the upper bound of privacy leakage based on the specific form of $\text{TV}(P^{\mathcal{O}}||P^{\mathcal{D}})$ of $\mathcal{M}$ according to this equation: $\widetilde{\epsilon}_{p}=2C_{1}-C_{2}\cdot{\text{TV}}(P^{\mathcal{O}}||P^{\mathcal{D}})$. In this section, we elaborate on the methods we propose to estimate $C_{1}$ and $C_{2}$. First, we introduce the following assumption.
Assumption E.1.
@@ Appendix E Practical Estimation for Average Bayesian Privacy
E.1 Estimation for $C_{1}=\sqrt{{\text{JS}}(F^{\mathcal{O}}||F^{\mathcal{B}})}$
@@
E.2 Estimation for $C_{2}=\frac{1}{2}(e^{2\xi}-1)$
Recall that the Maximum Bayesian Privacy, is defined as $\xi=\max_{w\in\mathcal{W},d\in\mathcal{D}}\left|\log\left(\frac{f_{D|W}(d|w)}{f_{D}(d)}\right)\right|$ and represents the maximal privacy leakage across all possible information $w$ released. The quantity $\xi$ remains constant when the extent of the attack is fixed. Under Assumption E.1, we have that
@@ E.2 Estimation for $C_{2}=\frac{1}{2}(e^{2\xi}-1)$
Ultimately, the computation of $C_{1}$ and $C_{2}$ hinges on the estimation of $\hat{f}_{{D}|{W}}(d|w_{m})$, which is further elaborated upon in Section D.1.
E.3 Estimation Error of ABP
In this section, we analyze the estimation error of ABP. To facilitate the analysis of the error bound, we consider that both the data and the model parameter are in discrete form. For the detailed analysis, please refer to [64].
@@ E.3 Estimation Error of ABP
Theorem E.9 (The Estimation Error of $C_{1}$).
@@
Appendix F Analysis for Theorem 5.5
In this section, we introduce the relationship between privacy leakage and distortion extent.
Now we introduce the definition of privacy leakage. The privacy leakage is measured using the gap between the estimated dataset and the original dataset. The semi-honest attacker uses an optimization algorithm ([67, 29, 66, 57]) to reconstruct the original dataset of the client given the exposed model information. Let $d^{(i)}_{m}$ represent the reconstructed $m$-th data at iteration $i$. Let $D$ be a positive constant satisfying that $||d^{(i)}_{m}-d||\in[0,D]$, $\forall i$ and $m$.
The privacy leakage is measured using protection extent and attacking extent. Assume that the semi-honest attacker uses an optimization algorithm to infer the original dataset of client $k$ based on the released parameter $w$. Let $\Delta=||W-\breve{W}||$ represent the distortion of the model parameter, where $\breve{W}$ represents the original parameter, and $W$ represents the protected parameter. The regret of the optimization algorithm in a total of $I$ rounds is $\Theta(I^{p})$.
Definition F.1 (Privacy Leakage).
Let $d$ represent the original private dataset, and $d^{(i)}$ represent the dataset inferred by the attacker at iteration $i$, $I$ represent the total number of learning rounds. The privacy leakage $\epsilon_{p}$ is defined as
$\epsilon_{p}=\left\{\begin{array}[]{cl}\frac{D-\frac{1}{I}\sum_{t=1}^{T}\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{I}^{(m)}-\breve{d}^{(m)}||}{D},&T>0\\ 0,&T=0\\ \end{array}\right.$ (36)
Remark:
(1) We assume that $||d^{(i)}-d||\in[0,D]$. Therefore, $\epsilon_{p}\in[0,1]$.
(2) If the estimated data of the adversary $d^{(i)}$ is equal to the original data $d$, i.e., $d^{(i)}=d$, then the privacy leakage is maximal. In this case, the privacy leakage is equal to 111.
(3) When the adversary does not attack ($I=0$), the privacy leakage $\epsilon_{p}=0$ from Eq. (36).
To derive bounds for privacy leakage, we need the following assumptions.
Assumption F.1.
Assume that $||d||\leq 1$.
Assumption F.2.
For any two datasets $d_{1}$ and $d_{2}$, assume that $c_{a}||g(d_{1})-g(d_{2})||\leq||d_{1}-d_{2}||\leq c_{b}||g(d_{1})-g(d_{2})||$.
Assumption F.3.
Assume that $c_{0}\cdot I^{p}\leq\sum_{i=1}^{I}||g(d^{(i)})-g(d)||=\Theta(I^{p})\leq c_{2}\cdot I^{p}$, where $d^{(i)}$ represents the dataset reconstructed by the attacker at round $I$, $d$ represents the dataset satisfying that $g(d)=w$, and $g(d^{(i)})=\frac{\partial\mathcal{L}(d^{(i)},w)}{\partial w}$ represents the gradient of the reconstructed dataset at round $I$.
With the regret bounds of the optimization algorithms, we are now ready to provide bounds for privacy leakage. The following lemma presents bounds for privacy leakage, following the analyses in [60, 62, 65].
Lemma F.11.
Assume that the semi-honest attacker uses an optimization algorithm to infer the original dataset of client $k$ based on the released parameter $W$. Let $d$ represent the dataset satisfying that $g(d)=W$, and $\breve{d}$ represent the dataset satisfying that $g(\breve{d})=\breve{W}$, where $\breve{W}$ represents the original parameter, $W$ represents the protected parameter, and $g(d)=\frac{\partial\mathcal{L}(W,d)}{\partial W}$. Let $d^{(m)}$ represent $m$-th data of dataset $d$, $\breve{d}^{(m)}$ represent $m$-th data of dataset $\breve{d}$. Let $\Delta=\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d^{(m)})-g(\breve{d}^{(m)})||$ represent the distortion of the parameter. The expected regret of the optimization algorithm in a total of $I$ ($I>0$) rounds is $\Theta(I^{p})$. If $\Delta\geq\frac{2c_{2}c_{b}}{c_{a}}\cdot I^{p-1}$, then we have that
$\epsilon_{p}\leq 1-\frac{c_{a}\cdot\Delta+c_{a}\cdot c_{0}\cdot I^{p-1}}{4D}.$ (37)
Proof.
Recall the privacy leakage of client $k$ $\epsilon_{p}^{(k)}$ is defined as
$\epsilon_{p}^{(k)}=\left\{\begin{array}[]{cl}\frac{D-\frac{1}{I}\sum_{i=1}^{I}\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{i}^{(m)}-\breve{d}^{(m)}||}{D},&I>0\\ 0,&I=0\\ \end{array}\right.$ (38)
To protect privacy, client $k$ selects a protection mechanism $M_{k}$, which maps the original parameter $\breve{W}$ to a protected parameter $W$. After observing the protected parameter, a semi-honest adversary infers the private information using the optimization approaches. Let $d_{i}$ represent the reconstructed data at iteration $i$ using the optimization algorithm. Therefore the cumulative regret over $I$ rounds
$R(I)$ $=\sum_{i=1}^{I}[||g(d_{i}^{(m)})-W^{(m)}||-||g(d^{(m)})-W^{(m)}||]$
$=\sum_{i=1}^{I}[||g(d_{i}^{(m)})-g(d^{(m)})||]$
$=\Theta(I^{p}).$
Therefore, we have
$c_{0}\cdot I^{p}\leq\sum_{i=1}^{I}||g(d_{i}^{(m)})-g(d^{(m)})||=\Theta(I^{p})\leq c_{2}\cdot I^{p},$ (39)
where $c_{0}$ and $c_{2}$ are constants independent of $I$.
Let $x$ and $\widetilde{x}$ represent two data. From our assumption, we have that
$c_{a}||g(x)-g(\widetilde{x})||\leq||x-\widetilde{x}||\leq c_{b}||g(x)-g(\widetilde{x})||.$ (40)
Let $d^{(m)}_{i}$ represent the reconstructed $m$-th data at iteration $i$ using the optimization algorithm. We have that
$\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{i}^{(m)}-\breve{d}^{(m)}||$ $\geq\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d^{(m)}-\breve{d}^{(m)}||-\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{i}^{(m)}-d^{(m)}||$
$\geq c_{a}\cdot\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d^{(m)})-g(\breve{d}^{(m)})||-c_{b}\cdot\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d_{i}^{(m)})-g(d^{(m)})||$
$\geq c_{a}\cdot||\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(d^{(m)})-g(\breve{d}^{(m)}))||-c_{b}\cdot\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d_{i}^{(m)})-g(d^{(m)})||$
where the second inequality is due to $\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d^{(m)}-\breve{d}^{(m)}||\geq c_{a}\cdot\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d^{(m)})-g(\breve{d}^{(m)})||$ and $\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{i}^{(m)}-d^{(m)}||\leq\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}c_{b}||g(d_{i}^{(m)})-g(d^{(m)})||$.
We denote $g^{(k)}$ as the original gradient of client $k$, and $\widetilde{g}^{(k)}$ as the distorted gradient uploaded from client $k$ to the server. Therefore, we have that
$\|\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(d^{(m)})-g(\breve{d}^{(m)}))\|$ $=\|\widetilde{g}^{(k)}-g^{(k)}\|$ (41)
$=\|\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(\breve{d}^{(m)})+\delta^{(k)}-g(\breve{d}^{(m)}))\|$ (42)
$=\|\delta^{(k)}\|$ (43)
$=\Delta^{(k)}.$ (44)
Therefore, we have
$D(1-\epsilon_{p}^{(k)})=\frac{1}{I}\sum_{i=1}^{I}\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||d_{i}^{(m)}-\breve{d}^{(m)}||$ $\geq c_{a}\Delta^{(k)}-c_{b}\cdot\frac{1}{I}\cdot\frac{1}{|\mathcal{D}^{(k)}|}\sum_{i=1}^{I}\sum_{m=1}^{|\mathcal{D}^{(k)}|}||g(d_{i}^{(m)})-g(d^{(m)})||$
$\geq c_{a}\Delta^{(k)}-c_{2}\cdot c_{b}I^{p-1}$
$\geq\frac{1}{2}\max\{c_{a}\Delta^{(k)},c_{2}\cdot c_{b}I^{p-1}\}$
$\geq\frac{c_{a}\Delta^{(k)}+c_{2}\cdot c_{b}I^{p-1}}{4}.$
Therefore, we have that
$\epsilon_{p}^{(k)}\leq 1-\frac{c_{a}\Delta^{(k)}+c_{2}\cdot c_{b}I^{p-1}}{4D}.$
Therefore, we have that
$\epsilon_{p}^{(k)}$ $\leq 1-\frac{c_{a}\Delta^{(k)}+c_{2}\cdot c_{b}I^{p-1}}{4D}$ (45)
$=1-\frac{c_{a}\|\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(d^{(m)})-g(\breve{d}^{(m)}))\|+c_{2}\cdot c_{b}I^{p-1}}{4D}.$ (46)
Appendix G Analysis for Theorem 5.5
Theorem G.12 (Relationship between Privacy and Robustness).
Let the privacy leakage of the protection mechanism $\mathcal{A}$ be measured using Definition 5.1, and let the robustness of $\mathcal{A}$ be measured using Definition 5.2. Let the privacy leakage be $\epsilon_{p}$, then $\left(r,\frac{Cr}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}\right)$, where $c_{a}$ and $D$ represent constants, $\mathcal{D}$ represents the dataset, and $\gamma$ represents the error rate.
Proof.
The distortion extent measures the discrepancy between the gradient of the original data and the distorted data. From the relationship between distortion extent and privacy leakage (Lemma 5.4), we know that the distortion extent is bounded as follows:
$\|\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(d^{(m)})-g(\breve{d}^{(m)}))\|\leq\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{c_{a}}.$ (47)
On one hand, we have that
$\|\nabla\mathcal{L}(\theta,(X+\delta,y))-\nabla\mathcal{L}(\theta,(X,y))\|$ $=\|\frac{1}{|\mathcal{D}^{(k)}|}\sum_{m=1}^{|\mathcal{D}^{(k)}|}(g(d^{(m)})-g(\breve{d}^{(m)}))\|$ (48)
$\leq\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{c_{a}}.$ (49)
On the other hand, from Lipschitz’s condition, we know that
$\|\nabla\mathcal{L}(\theta,(X+\delta,y))-\nabla\mathcal{L}(\theta,(X,y))\|\leq C\|\delta\|,$ (50)
where $C$ represents a constant. Therefore, we have that
$\|\nabla\mathcal{L}(\theta,(X+\delta,y))-\nabla\mathcal{L}(\theta,(X,y))\|\leq\frac{C\|\delta\|}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}.$ (51)
Therefore, we have that
$\mathbb{E}_{X\sim P}[\text{sup}_{\|\delta\|_{p}\leq r}|\nabla\mathcal{L}(\theta,(X+\delta,y))-\nabla\mathcal{L}(\theta,(X,y))|]$ (52)
$\leq\frac{C\|\delta\|}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}$ (53)
$\leq\frac{Cr}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}.$ (54)
Recall that a model is $(r,\epsilon)$-input-robust, if it holds that
$\mathbb{E}_{X\sim P}[\text{sup}_{\|\delta\|_{p}\leq r}|\nabla\mathcal{L}\left(w,(X+\delta,y)\right)-\nabla\mathcal{L}\left(w,(X,y)\right)|]\leq\epsilon,$ (55)
where $X$ denotes the data (training sample or feature), and $y$ denotes the corresponding label.
Therefore, from the definition of robustness, we know that the model is $\left(r,\frac{Cr}{2}+\frac{4D(1-\epsilon_{p})-c_{2}c_{b}I^{p-1}}{2c_{a}}\right)$-input-robust. ∎